Hours per Unit Production 0.60 22,000 units 0.20 43,000 units Annual Hubs Sprockets Additional information about the company follows a. Hubs require $21 in direct materials per unit, and Sprockets require $17 b. The direct labor wage rate is 13 per hour c. Hubs are more complex to manufacture than Sprockets and they require special equipment. d. The ABC system has the following activity cost pools Activity Estimated Activity Cost Pool (Activity Measure) Machine setups (number of setups) Special processing (machine-hours) General factory (organization-sustaining) Overhead Cost Hubs 150 3,900 NA Sprockets Total 120 $ 32,400 $195,000 $185,200 270 0 3,900 NA NA Required: 1. Compute the activity rate for each activity cost pool ctivity Cost Machine setups Special processing Activity Rate per setu per MH 2. Determine the unit product cost of each product according to the ABC system. (Round intermediate calculations and final answers to 2 decimal places.) Hubs Sprockets Direct materials Direct labor Overhead Unit costFogerty Company makes two products, titanium Hubs and Sprockets. Data regarding the two products follow:
Answer #1
1) Calculation of activity rate for each cost pool
Activity cost pools Estimated overhead(a) Estimated cost driver(b) activity rates((a)/(b))
Machine setup 32400 270 setups 120.00
Special Processing 195000 3900 machine hours 50.00
Machine setup= $120 per setup
Special processing=$50 per MH
2) Calculation of unit product cost:
Particulars Hubs Sprockets
Direct material 21 17
Direct labour( hubs=13*0.6, sprockets= 13*0.2) 7.8 2.6
Overhead 9.68 0.33
Unit product cost 38.48 19.93
Working:
Calculation of total overhead of hubs:
Machine setups (150*120)= 18000
Special processing=3900*50=195000
Total overhead =18000+195000= 213000
Overhead cost per unit=213000/22000=$9.68
Calculation of total overhead of sprokets:
Machine setups (120*120)= 14400
Total overhead =14400
Overhead cost per unit=14400/43000= $0.33